# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v3

@@ 3.3 Identification of the shocks in structural TVAR models
Suppose the transition weights are binary, $\alpha_{m,t}\in\{0,1\}$, for all $t$ and $m=1,...,M$. Then, at every $t$, the process is completely in one of the regimes and the impact matrix is $B_{y,t}=B_{m}$ for the regime $m$ with $\alpha_{m,t}=1$. That is, we have a structural TVAR model. By Lemma 2, the impact matrix $B_{y,t}$ is thereby identified up to ordering and signs of its columns in each of the regimes, implying that identification is obtained with constraints that fix the ordering and signs of the columns the impact matrix in each regime. To state the identification result, a specific shock distribution is assumed here for simplicity. Namely, we assume that each shock $e_{it}$, $i=1,...,d$, follows the skewed $t$-distribution of Hansen (1994), which can flexibly capture fat tails and skewness in the data, nests the Gaussian distribution as a special case, and is discussed in detail in Section 4.111The identification result can be extended to other suitable shock distributions using arguments similar to those in the proof of Proposition 1 in Appendix B.3.
The following proposition, which is proved in Appendix B.3, states the identification result for the TVAR models.
@@
Consider the STVAR model defined in Equations (3.2) and (3.3) with $B_{1},...,B_{M}$ invertible, Assumption 1 satisfied, $\alpha_{m,t}\in\{0,1\}$ for all $t$ and $m=1,...,M$, and the shocks $e_{it}$, $i=1,...,d$, following the skewed $t$-distribution defined in Equation (4.1) with the parameters $\nu_{i}>2$ and $\lambda_{i}\in(-1,1)$ controlling kurtosis and skewness, respectively. Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
@@
Then, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
Condition (a) of Proposition 1 fixes the ordering and signs of $B_{1}$, which effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ under Conditions (b) and (c), respectively. Any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality (e.g., by assuming that the first nonzero entry in each column is positive and the first nonzero entries are in a decreasing order, given that none of them are equal). If Condition (b) fails for some $i,j\in\{1,...,d\}$, the distributions of these shocks are identical, implying that fixing the ordering of the columns of $B_{1}$ does not fix the ordering corresponding of the columns of $B_{2},...,B_{M}$. In such a case, the ordering of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality, yielding identification. If Condition (c), in turn, fails for some $i\in\{1,...,d\}$, the signs of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality, yielding identification. The result of Proposition 1 holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$, because in some special cases (e.g., if all observations are zero), changing the ordering of the columns of $B_{2},...,B_{M}$ does not change the joint density of the model.
If Condition (b) or (c) is close to break, the identification is weak with respect to the ordering or signs of the corresponding columns of $B_{2},...,B_{M}$, respectively. Somewhat surprisingly, our experience shows that even when Conditions (b) and (c) are not close to break, the identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, at least if the parameters are estimated by the method of maximum likelihood or penalized maximum likelihood (discussed in Section 4). Specifically, there are frequently a number of local maxima of the (penalized) log-likelihood function with (penalized) log-likelihoods close to each other, corresponding to otherwise largely similar parameter estimates but with different the ordering and signs of the columns of $B_{2},...,B_{M}$. As discussed in Section 3.5, this weak identification should be appropriately dealt with when labelling the shocks, for instance, by combining Proposition 1 with additional identifying information.
@@ 3.4 Identification of the shocks in structural STVAR models
Having established the identification result for structural TVAR models, we now consider the transition weights that exhibit more variation. The transition weights are here assumed to be either exogenous (nonrandom) or logistic as defined in (2.4), so that the result of Lemma 2 establishing identification of the transition weights can be exploited. Because the variation in the transition weights translates to variation in the impact matrix (3.3), which is at each $t$ identified up to ordering and signs of its columns by Lemma 2, the identification result can be stated without making further assumptions about the distributions of the shocks other than those in Assumption 1. Specifically, under the conditions of the following proposition (which is proven in Appendix B.4), the matrices $B_{1},...,B_{M}$ are uniquely identified.
@@
Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 satisfied, and let the transition weights be either exogenous (nonrandom) or follow the logistic process defined in (2.4) (with $M=2$ assumed). Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
@@
    the ordering and signs of the columns of $B_{1}$ are fixed,
@@
    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero, and
  3. (C)
    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors take at least $M+1$ linearly independent values.
Then, for models with logistic transition weights, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$, and $z_{t}\in\mathbb{R}^{T}$, where $z_{t}$ is the switching variable defined in (2.4). For models with nonrandom exogenous transition weights, this result holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
Proposition 2 essentially states that if the structural shocks are independent and at most one of them is Gaussian (Assumption 1), the impact matrices $B_{1},...,B_{M}$, and hence, the structural shocks are identified. The result holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, as there may exist some special cases of $B_{1},...,B_{M}$ such that fixing the ordering and signs of the columns of $B_{1}$ does not fix the ordering and signs of the columns of $B_{2},...,B_{M}$, but this set has a Lebesgue measure zero in $\mathbb{R}^{d\times dM}$. For logistic models, the identification holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$, as special cases of degenerate data may lack sufficient variation to uniquely identify the transition weights (see Lemma 2).
Condition (A) of Proposition 2 states that the ordering and signs of the columns of $B_{1}$ should be fixed, which fixes the ordering and signs of the columns of $B_{y,t}$. If the shocks follow the skewed $t$-distributions defined in (4.1), any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality, as reordering the columns would just reorder the shocks and swapping a sign of a column corresponds to swapping the sign of the related skewness parameter. Condition (B) states that the transition weights should, for some $t$, be strictly positive for all of the regimes. If there are only two regimes, the requirement is merely that the regime-switches are not always discrete, in which case Proposition 1 would be more relevant.
Condition (C) of Proposition 2 states that the transition weights should exhibit at least a certain small degree of variation. For instance, if there are only two regimes in the model, the transition weights should take at least $M+1$ distinct values, as the vectors $(\alpha_{1,t},\alpha_{2,t})$ with distinct $\alpha_{1,t}$ are always linearly independent due to the constraint $\sum_{m=1}^{M}\alpha_{m,t}=1$. This assumption is not, however, very restrictive, as the transition weights vary either depending on the preceding observations or some exogenous variables, and it is very uncommon that their variation would result in only a few distinct values of the transition weights (except when the regime switches are discrete). Similarly to the structural TVAR model discussed in Section 3.3, our experience shows that the identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately dealt with when labelling the shocks, as discussed in the next section.
@@ 3.5 Labelling the shocks and blended identification
As in the linear SVAR model of Lanne et al. (2017), the statistically identified structural shocks do not necessarily have an economic interpretation, and labelling them as economic shocks requires external information. Our identification method, nonetheless, facilitates identifying structural shocks with time-varying impact effects without imposing restrictions on the impact responses of the variables. This is a major advantage compared to statistical identification by heteroscedasticity that restricts how the impact responses may vary over time (see Lütkepohl and Netšunajev, 2017). Additionally, in line with the statistical identification literature, many economically motivated overidentifying restrictions are testable, as the shocks are statistically identified.
Labelling the shocks based on the estimates of the impact matrices $B_{1},...,B_{M}$ might not always be straightforward, as the same shock needs to be associated with the same column of the impact matrix $B_{m}$ in all regimes. For example, if a positive supply shock increases output and decreases prices on impact, and it should do so in all regimes, labelling the $i$th shock as the supply shock requires the $i$th column of all $B_{1},...,B_{M}$ to satisfy such signs. However, as discussed in Sections 3.3 and 3.4, the identification is often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately accounted for when labelling the shocks.
Assuming that the parameters are estimated by the method of maximum likelihood (ML) or penalized maximum likelihood (PML) described in Section 4, weak identification often results in multiple local maxima of the (penalized) log-likelihood function with the related (penalized) log-likelihoods close to each other. If the identification is weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, these local solutions often correspond to otherwise largely similar parameter estimates but with different ordering and signs of the columns of $B_{2},...,B_{M}$. As our estimation method, discussed in Section 4, can produce a set of local solutions, the possible presence of such local solutions can be easily checked.
@@ 4.1 Maximum likelihood estimation and ergodic stationarity
Maximizing the log-likelihood function (presented in Section 4.2) can be challenging in practice due to its high multimodality, induced by the nonlinear dynamics. Imposing the stability condition can make estimation particularly difficult when the data is persistent, as in such cases numerical optimization algorithms frequently gravitate to the boundary of the parameter space, where they perform poorly and are not often even able to converge to a local maximum. To overcome this issue, we propose allowing for unstable estimates, but adding a penalty term to the log-likelihood function that penalizes parameter values falling outside or close the boundary of the stability region, and then maximizing the penalized log-likelihood function. In this way, the optimization algorithm can explore the parameter space also outside the stability region, facilitating improved performance of the optimization algorithms, while sufficient penalization eventually steers the optimization algorithm back to the stability region.
@@ 4.2 The penalized log-likelihood function
where $\mathbbm{1}\{e_{it}<-a_{i}/b_{i}\}$ is an indicator function that takes the value one if $e_{it}<-a_{i}/b_{i}$ and zero otherwise, and $\mathbbm{1}\{e_{it}\geq-a_{i}/b_{i}\}$ is an indicator function that takes the value one if $e_{it}\geq-a_{i}/b_{i}$ and zero otherwise. The constants $a_{i},b_{i}$, and $c_{i}$ are defined as $a_{i}=4\lambda_{i}c_{i}\left(\frac{\nu_{i}-2}{\nu_{i}-1}\right)$, $b_{i}=(1+3\lambda_{i}^{2}-a_{i}^{2})^{1/2}$, and $c_{i}=\frac{\Gamma\left(\frac{\nu_{i}+1}{2}\right)}{(\pi(\nu_{i}-2))^{1/2}\Gamma\left(\frac{\nu_{i}}{2}\right)}$, where $\Gamma(\cdot)$ is Gamma function.
@@
where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix, $\mu_{y,t}$ is defined in Section 2 and $B_{y,t}$ in Equation (3.3). The parameters of the model are collected to the vector $\boldsymbol{\theta}=(\phi_{1,0},...,\phi_{M,0},$ $\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))$, $m=1,...,M$, $\sigma=(\text{vec}(B_{1}),...,\text{vec}(B_{M}))$, $\alpha$ contains the parameters governing the transition weights, and $\nu=(\nu_{1},...,\nu_{d},\lambda_{1},...,\lambda_{d})$ contains the degrees-of-freedom and skewness parameters. [^2] With threshold (2.3) or logistic (2.4) weights, $\alpha=(r_{1},...,r_{M-1})$ or $\alpha=(c,\gamma)$, respectively, whereas with exogenous weights $\alpha$ is omitted from the parameter vector.
@@
where $L(\boldsymbol{\theta})$ is defined in (4.2) and $P(\boldsymbol{\theta})\geq 0$ is a penalization term that penalizes the log-likelihood function from parameter values that are close to entering or are in an uninteresting region of the parameter space. Focusing on avoiding unstable estimates, we define the penalization term as
@@
Whenever the companion form AR matrix of a regime has eigenvalues greater than $1-\eta$ in modulus, the penalization term (4.4) is greater than zero, and it increases in the modulus of these eigenvalues. Our penalization term incorporates the multiplicative coefficient $Td$ in attempt to standardize the strength of the penalization with respect to the number of observations $T$ and variables $d$, as the value of the log-likelihood function (4.2) tends to scale with these quantities. In our empirical application (Section 5), we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$, thus, penalizing parameter values outside the stability region significantly, while maintaining flexibility near the boundary.
4.3 Three-phase estimation procedure
Lanne et al. (2017) propose a three-phase estimation procedure for their linear SVAR model identified by non-Gaussianity. In the first phase, least squares (LS) estimation produces preliminary estimates of certain parameters, which serve as initial values in numerical optimization algorithms used in the subsequent phases to maximize the log-likelihood function. The nonlinear dynamics of our structural STVAR model, however, add substantial challenges to the estimation problem by inducing a high number of modes to the (penalized) log-likelihood function. To address these challenges, we propose a modified version of the three-phase estimation procedure in which the multimodality is explicitly taken into account.
Our three-phase estimation procedure operates with the following steps:

### v4

@@ 3.3 Identification of the shocks in structural TVAR models
Suppose the transition weights are binary, $\alpha_{m,t}\in\{0,1\}$, for all $t$ and $m=1,...,M$, specifically being either exogenous or of the threshold form (2.3). Then at every $t$, the process is completely in one of the regimes, making the impact matrix $B_{y,t}=B_{m}$ for the active regime $m$ with $\alpha_{m,t}=1$. By Lemma 2, $B_{y,t}$ is thereby identified up to ordering and signs of its columns within each regime, and identification requires constraints that fix these orderings and signs for each regime-specific impact matrix $B_{1},...,B_{M}$. For simplicity, and to state the identification result under just-identifying constraints, we assume each shock $e_{it}$, $i=1,...,d$, follows the skewed $t$-distribution of Hansen (1994), which can flexibly capture fat tails and skewness, and nests the Gaussian distribution as a special case. The skewed $t$-distribution is described in Section 4 discussing estimation of the model parameters. [^1] The identification result can be extended to other suitable shock distributions using arguments similar to those in the proof of Proposition 1 in Appendix B.3.
The following proposition, which is proved in Appendix B.3, states the identification result for the TVAR model and models with exogenous binary weights.
@@
Consider the STVAR model defined in Equations (3.2) and (3.3) with $B_{1},...,B_{M}$ invertible; Assumption 1 and the conditions of Lemma 2 satisfied; $\alpha_{m,t}$ either of the threshold form (2.3) or exogenous (nonrandom) with $\alpha_{m,t}\in\{0,1\}$ for all $t$ and $m=1,...,M$; and the shocks $e_{it}$, $i=1,...,d$, following the skewed $t$-distribution defined in Equation (4.1) with the parameters $\nu_{i}>2$ and $\lambda_{i}\in(-1,1)$ controlling kurtosis and skewness, respectively. Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
@@
Then, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T})\in\mathbb{R}^{d(T+p)}$.
Condition (a) of Proposition 1 fixes the ordering and signs of $B_{1}$, which effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ under Conditions (b) and (c), respectively. Any fixed ordering and signs of the columns of $B_{1}$ can be chosen without loss of generality (e.g., by making the first nonzero entry in each column is positive and ordering these entries in decreasing magnitude, given that none of them are equal). If Condition (b) fails for some $i,j\in\{1,...,d\}$, the distributions of these shocks are identical, implying that fixing the ordering of the columns of $B_{1}$ does not fix the ordering of the corresponding columns of $B_{2},...,B_{M}$. In such a case, the ordering of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality. Similarly, if Condition (c) fails for some $i\in\{1,...,d\}$ (meaning zero skewness), the signs of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality. The result of Proposition 1 holds only almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T})\in\mathbb{R}^{d(T+p)}$, because in some special cases that comprise set with a Lebesgue measure zero, changing the ordering of the columns of $B_{2},...,B_{M}$ does not change the joint density of the model.
If Condition (b) or (c) is close to break, identification is weak with respect to the ordering or signs of the corresponding columns of $B_{2},...,B_{M}$, respectively. Somewhat surprisingly, our experience shows that even when Conditions (b) and (c) hold clearly, identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, at least if the parameters are estimated by the method of maximum likelihood or penalized maximum likelihood (discussed in Section 4). Specifically, there are frequently a number of local maxima of the (penalized) log-likelihood function with (penalized) log-likelihoods close to each other, corresponding to otherwise largely similar parameter estimates but with different the ordering and signs of the columns of $B_{2},...,B_{M}$. As discussed in Section 3.5, this weak identification should be appropriately dealt with when labelling the shocks, for instance, by combining Proposition 1 with additional identifying information.
@@ 3.4 Identification of the shocks in structural STVAR models
Having established the identification result for structural TVAR models, we now consider transition weights that exhibit more variation. In particular, the transition weights are here assumed to be either exogenous (nonrandom) or logistic as defined in (2.4), so that the result of Lemma 2 establishing identification of the transition weights can be exploited. Under the conditions of the following proposition (which is proven in Appendix B.4), the matrices $B_{1},...,B_{M}$ are uniquely identified.
@@
Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 and the conditions of Lemma 2 satisfied, and let the transition weights be either exogenous (nonrandom) or follow the logistic process defined in (2.4) (with $M=2$ assumed). Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
@@
    the ordering and signs of the columns of $B_{1}$ are fixed, and
@@
    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero.
Then, for models with logistic transition weights, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$. For models with nonrandom exogenous transition weights, this result holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
Proposition 2 essentially states that if the structural shocks are independent and at most one of them is Gaussian (Assumption 1), the impact matrices $B_{1},...,B_{M}$, and hence, the structural shocks are identified. The result holds only almost everywhere, because the parameters are not identified in some special cases that constitute a set with Lebesgue measure zero (see also Lemma 2).
Condition (A) of Proposition 2 states that the ordering and signs of the columns of $B_{1}$ should be fixed, which fixes the ordering and signs of the columns of $B_{y,t}$. If the shocks follow the skewed $t$-distributions defined in (4.1), any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality, as reordering the columns would just reorder the shocks and swapping a sign of a column corresponds to swapping the sign of the related shock and skewness parameter. Condition (B) states that the transition weights should, for some $t$, be strictly positive for all of the regimes. If there are only two regimes, the requirement is merely that the regime-switches are not always discrete, in which case Proposition 1 would be relevant. Similarly to the structural TVAR model discussed in Section 3.3, our experience shows that identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately dealt with when labelling the shocks, as discussed in the next section.
@@ 3.5 Labelling the shocks and blended identification
As in the linear SVAR model of Lanne et al. (2017), the statistically identified structural shocks do not necessarily have an economic interpretation, and labelling them as economic shocks requires external information. Our identification method, nonetheless, facilitates identifying structural shocks with time-varying impact effects without imposing restrictions on the impact responses of the variables. This is a major advantage compared to statistical identification by heteroscedasticity that restricts how the impact responses may vary over time (see Lütkepohl and Netšunajev, 2017).
Labelling the shocks based on the estimates of the impact matrices $B_{1},...,B_{M}$ might not always be straightforward, as the same shock needs to be associated with the same column of the impact matrix $B_{m}$ in all regimes. For example, if a positive supply shock increases output and decreases prices on impact, and it should do so in all regimes, labelling the $i$th shock as the supply shock requires the $i$th column of all $B_{1},...,B_{M}$ to satisfy such signs. However, as discussed in Sections 3.3 and 3.4, identification can be weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately accounted for when labelling the shocks.
Assuming that the parameters are estimated by the method of maximum likelihood (ML) or penalized maximum likelihood (PML) described in Section 4, weak identification often results in multiple local maxima of the (penalized) log-likelihood function with the related (penalized) log-likelihoods close to each other. If identification is weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, these local solutions often correspond to otherwise largely similar parameter estimates but with different ordering and signs of the columns of $B_{2},...,B_{M}$.
@@ 4.1 Maximum likelihood estimation and ergodic stationarity
Maximizing the log-likelihood function (presented in Section 4.2) can be challenging in practice due to its high multimodality, induced by the nonlinear dynamics. Imposing the stability condition can make estimation particularly difficult when the data is persistent, as in such cases numerical optimization algorithms frequently gravitate to the boundary of the parameter space, where they perform poorly and are not often even able to converge to a local maximum. To overcome this issue, we propose to allow for unstable estimates and to maximize the penalized log-likelihood function, obtained by adding a penalty term to the log-likelihood function that penalizes parameter values falling outside or close to the boundary of the stability region. In this way, the optimization algorithm can explore the parameter space also outside the stability region, facilitating improved performance, while sufficient penalization eventually steers the optimization algorithm back to the stability region.
@@ 4.2 The penalized log-likelihood function
where $\mathbbm{1}\{e_{it}<-a_{i}/b_{i}\}$ is an indicator function that takes the value one if $e_{it}<-a_{i}/b_{i}$ and zero otherwise, and $\mathbbm{1}\{e_{it}\geq-a_{i}/b_{i}\}$ is an indicator function that takes the value one if $e_{it}\geq-a_{i}/b_{i}$ and zero otherwise. The constants $a_{i},b_{i}$, and $c_{i}$ are defined as $a_{i}=4\lambda_{i}c_{i}\left(\frac{\nu_{i}-2}{\nu_{i}-1}\right)$, $b_{i}=(1+3\lambda_{i}^{2}-a_{i}^{2})^{1/2}$, and $c_{i}=\frac{\Gamma\left(\frac{\nu_{i}+1}{2}\right)}{(\pi(\nu_{i}-2))^{1/2}\Gamma\left(\frac{\nu_{i}}{2}\right)}$, where $\Gamma(\cdot)$ is the Gamma function.
@@
where $I_{d,i}$ is the $i$th column of the $d$-dimensional identity matrix, $\mu_{y,t}$ is defined in Section 2 and $B_{y,t}$ in Equation (3.3). The parameters of the model are collected to the vector $\boldsymbol{\theta}=(\phi_{1},...,\phi_{M},$ $\varphi_{1},...,\varphi_{M},\sigma,\alpha,\nu)$, where $\varphi_{m}=(\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))$, $m=1,...,M$, $\sigma=(\text{vec}(B_{1}),...,\text{vec}(B_{M}))$, $\alpha$ contains the parameters governing the transition weights, and $\nu=(\nu_{1},...,\nu_{d},\lambda_{1},...,\lambda_{d})$ contains the degree-of-freedom and skewness parameters. [^2] With threshold (2.3) or logistic (2.4) weights, $\alpha=(r_{1},...,r_{M-1})$ or $\alpha=(c,\gamma)$, respectively, whereas with exogenous weights, $\alpha$ is omitted from the parameter vector.
@@
where $L(\boldsymbol{\theta})$ is defined in (4.2) and $P(\boldsymbol{\theta})\geq 0$ is a penalization term that penalizes the log-likelihood function from parameter values that are close to entering or are in an uninteresting region of the parameter space. To focus on avoiding unstable estimates, we define the penalization term as
@@
Whenever the companion form AR matrix of a regime has eigenvalues greater than $1-\eta$ in modulus, the penalization term (4.4) is greater than zero, and it increases in the modulus of these eigenvalues. Our penalization term incorporates the multiplicative coefficient $Td$ in an attempt to standardize the strength of penalization with respect to the number of observations $T$ and variables $d$, as the magnitude of the log-likelihood typically increases with these quantities. In our empirical application (Section 5), we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$, thus, penalizing parameter values outside the stability region significantly, while maintaining flexibility near the boundary.
4.3 Three-step estimation procedure
Lanne et al. (2017) propose a three-step estimation procedure for their linear SVAR model identified by non-Gaussianity. In the first step, least squares (LS) estimation produces preliminary estimates of certain parameters, which serve as initial values in numerical optimization algorithms used in the subsequent steps to maximize the log-likelihood function. The nonlinear dynamics of our structural STVAR model, however, add substantial challenges to the estimation problem by inducing a high number of modes to the (penalized) log-likelihood function. To address these challenges, we propose a modified version of the three-step estimation procedure in which multimodality is explicitly taken into account.
Our three-step estimation procedure proceeds with the following steps: